OpenStudy (anonymous):

Approximately how many orders of magnitude longer is one year than one second?

OpenStudy (anonymous):

1sec is 1sec 1min=60seconds 60mins=1hour 24hours=1day 365 days=1 year so 60 x 60x24 x365=? 1 order of magnitude is about 10 times larger or longer according to your problem 2 order of magnitude is about 100 times larger 3 order of magnitude is about 1000 times larger and so on Example 45672 is 4 magnitude of order higher than 3 but is 3 magnitude higher than 13 or 78

OpenStudy (michele_laino):

hint: seconds contained in one year are: \[\Large 365 \times 24 \times 60 \times 60 = {\text{31}}{\text{,536}}{\text{,000}} \cong {\text{3}} \cdot {\text{1}}{{\text{0}}^{\text{7}}}\]